OpenStudy (anonymous):

I WILL MEDAL AND FAN i will medal let a,b,c be points (2,3), (3,-2) and (-1,4) find lenght BC and AC 30−−√,42−−√,25−−√,52−− for BC please help me.

OpenStudy (anonymous):

√30,√42,√25,\[\sqrt{52}\] ARE OPTIONS for BC

OpenStudy (anonymous):

let a,b,c be points (2,3), (3,-2) and (-1,4) find lenght BC and AC

OpenStudy (anonymous):

ok so youre trying to find the length, alright ill try to help a little bit

OpenStudy (anonymous):

WHERE \[\sqrt{30},\sqrt{42},\sqrt{25},\sqrt{52} \] ARE THE BC options

OpenStudy (anonymous):

bc is \[\sqrt{52}\]